# From Simulations to Reality: Dark Energy Reconstruction with Simulated SNIa data from the Vera C. Rubin Observatory

### v2

@@ IV.2 Pantheon and Pantheon+ compilations
The Pantheon datasets are all publicly available in the form of unbinned Hubble diagram data. Pantheon has 1048 data points scolnic2018complete within redshifts between $z=0.01$ and $z=2.3$, and Pantheon+ has 1550 Type Ia between $z=0.001$ and $z=2.26$ brout2022pantheon+ . They are the latest compilations of the SNIa observations, expanding from the previously concluded confirmed Pan-STARRS1 survey 2002SPIE.4836..154K . Unlike the LSST simulated data, all of Pantheon’s data are spectroscopic in nature.
@@ V Methodology
This section describes how we use the ANN to generate neural network reconstructions for the distance modulus. We employ the methodology presented in gomez2023neuralrecs using the hyperparameter tuning suggested in gomez2023neuralgenetic . In our case, we use the TensorFlow library to implement our neural network model. Below we outline the steps:
@@
  • We prepare the dataset for the training of our neural network. We consider the redshift ($z$) as input, and the distance modulus ($\mu$) with the error ($\sigma(\mu)$) conformed by the sum of their statistical and systematic errors.
  • We explore the dataset and train different architectures of neural networks with Monte Carlo Dropout gal2015dropout (as in gomez2023neuralrecs ) to gain empirical knowledge about the range of hyperparameters. The Monte Carlo Dropout allows to obtain the uncertainties of the neural network predictions and to generate a more robust model based on the data. We use the Monte Carlo dropout implementation from AstroANN astronnleung2019deep ; astronn .
@@
  • To find a proper combination of hyperparameters, we set the search space for the use of nnogada (gomez2023neuralgenetic, ) and their implemented simple genetic algorithms with elitism from the DEAP library DEAP_JMLR2012 ; de2014deap (for details on genetic algorithms we recommend Ref. medel2023cosmological ). In our case, we consolidated the search space with the number of layers $\in[3,4]$, the number of neurons for hidden layers $\in[100,200]$, and batch size $\in[8,16,32,64]$.
@@
  • The best combination found was 4 layers with 200 neurons in each one and a batch size of 32 with a loss function value of 0.02528 (see Fig. 1).
  • We use the best combination to build our ANN network and apply the training data to it.
@@
  • Once the neural network is trained, we perform several predictions of the distance modulus to generate a non-parametric reconstruction based on the LSST simulated SNIa data.
@@
  • We test with the two dark energy models discussed previously. We compare the theoretical predictions and the reconstruction with the data, and we measure with the MSE metric to test which cases are in more agreement with the dataset. The values of the dark energy model parameters are previously fitted with Bayesian inference using the same data used for the neural reconstruction.
@@ VI.1 Bayesian Analysis
To have an insight into the cosmological models considered in this work constrained by the LSST data, we perform an initial Bayesian inference using nested sampling with the dynesty library speagle2020dynesty and the SimpleMC code aubourg2015 ; simplemc . We also do it with Pantheon and Pantheon+, to compare the results obtained with the LSST simulations. This provides us with a baseline reference, before the subsequent ANN analysis.
Table 1 includes the ranges of the priors used for the free parameters of both $\Lambda$CDM and CPL. In addition, it shows the results of posterior distribution sampling and the Bayes factor to compare models. According to Fig. 2 and Table 1, we can notice that for $\Lambda$CDM the $\Omega_{m}$ is considerably better constrained with the LSST data than the other two SNIa compilations; while the $h$ parameter is practically not affected between the three different datasets. On the other hand, for the CPL model, the parameter estimation with LSST data is consistent with Pantheon and Pantheon+, with a better constraint for the $w_{0}$ parameter.
In all the cases the $-2\log\mathcal{L}_{max}$ (Table 1) obtains a lower value for the LSST dataset, which implies that this data obtains a better fit for the cosmological models. In addition, something interesting is that using the LSST data, the Bayesian model comparison suggests that $\Lambda$CDM increases its advantage over CPL because according to Jeffrey’s scale, both Pantheon and Pantheon+ $\Lambda$CDM only have weak evidence, and for LSST data this scale turns to moderate. This disadvantage can also be analyzed in Fig. 3, where the EoS for CPL using LSST data puts the $w=-1$ line for $\Lambda$CDM almost out of $1-\sigma$, while for Pantheon and Pantheon+, the most probable value for $w$ is statistically closer to the $\Lambda$CDM prediction. However, considering only the values of $-2\log\mathcal{L}_{max}$, for the three datasets, CPL has always a better fit to the data than $\Lambda$CDM; therefore, the Bayesian evidence could be very sensitive to the extra-parameters of CPL and, in future work, it could be worth to incorporate other model comparison techniques because Bayesian evidence has received some criticisms in cosmological data analysis efstathiou2008limitations ; nesseris2004comparison ; koo2022bayesian . However, our results clearly show that the LSST data can give more advantages while investigating for model selection tests in the future (more so with LSST $10$ years SNIa data) though in the current analysis our results from Table 1, are insufficient and very weak to choose in favor of any model.
@@
Evidence - - Weak - Weak - Moderate
Table 1: Bayesian analysis to $\Lambda$CDM and CPL models using Pantheon, Pantheon+ and LSST data. We can notice the priors of their parameters, their parameter estimation, $-2\log\mathcal{L}_{max}$, the log-Bayesian evidence $\log Z$, the log-Bayes factor $\log B$ of CPL and $\Lambda$CDM for each dataset. Log-Bayes factor is defined as $\ln B_{ab}\equiv\ln Z_{a}-\ln Z_{b}$, and according to Jeffrey’s scale wasserman2000bayesian we can consider weak evidence if $10\leq|\Delta\log Z|<$ 1, moderate evidence if $31\leq|\Delta\log Z|<$ 3, strong evidence if $53\leq|\Delta\ln Z|<$ 5, and decisive evidence if $|\Delta\log\mathcal{Z}|\geq 5$, in favourof the preferred model; if $\Delta\log Z$ is negative it is in favourof the second term, and if it is positive the first term is favoured hee2016bayesian . In our results, $\Lambda$CDM is always the preferred model.
@@
Figure 2: One and two-dimensional posterior distributions from Bayesian sampling for the free parameters of $\Lambda$CDM (left) and CPL (right) using SNIa data from Pantheon, Pantheon+, and LSST simulations.
@@ VI.2 Neural Reconstruction
Finally, after careful training of a neural network model with the LSST simulated data, we generated the model-independent reconstruction for $\mu(z)$ as can be seen in Fig. 4. In Fig.4 we plot the distance modulus ($\mu$) from the datasets of LSST sim and Pantheon, Pantheon+, and then we overlay with the reconstructed distance modulus (black dashed lines) along with their associated uncertainties from the ANN predictions plus the modeled errors (yellow regions). We also plot the theoretically computed $\mu(z)$ from the two cosmological dark energy models of $\Lambda$CDM and CPL. From the left panel of Fig. 4 it can be seen that the reconstructed data points from the ANN model and the LSST observations (green points) are very close. Next, we computed the MSE for each of these three cases with the observed LSST dataset, shown in Table 2. From Table 2 we see that MSE for ANN is slightly better than $\Lambda$CDM or CPL. This implies that the artificial neural network (ANN) based non-parametric reconstruction model can indeed more closely replicate the data than the theoretical models, albeit by a small margin. The second row in Table 2 gives additional insight by comparing the MSE between the ANN reconstructed $\mu(z)$ and the corresponding $\Lambda$CDM or CPL based distances. It is worth noting that, due to the hyperparameter tuning of the ANN and the Monte Carlo dropout, we have a good model based on the data, which looks visually smooth and does not have any overfitting or underfitting in machine learning terms.
@@
In Fig. 5 we present residual plots for the distance modulus $\mu(z)$, computed as $\delta\mu\mid_{i}=\mu_{i}-\mu_{model}(z_{i})$ (where $\mu_{i}$ corresponds to the $i^{th}$ observation from the LSST sim data) for two cases, a) LSST sim data - $\Lambda$CDM theory and b) LSST sim data - ANN model reconstructed $\mu(z)$ (we compare the $\mu(z)$ residuals between the ANN model and the $\Lambda$CDM model only, as $\Lambda$CDM is seen to have bigger differences with the ANN reconstruction as seen in Table 2). It can be seen that there is an infinitesimal difference between the residuals from $\Lambda$CDM and ANN model. However, we can notice the sensitivity of the ANN model from outliers, which can make the error margins get bigger. We performed an additional $F$-test to see if there is any statistically significant difference between the distance modulus residuals from $\Lambda$CDM and ANN model  [^3] The $F$-test is used to compare the variances of two independent samples to test the hypothesis that they are equal. It is defined as the ratio of the variances, $F=\frac{\sigma_{1}^{2}}{\sigma_{2}^{2}}$, where $s_{1}^{2}$ and $s_{2}^{2}$ are the sample variances. This statistic follows the $F$-distribution under the null hypothesis that the variances are equal snedecor1989statistical . From the $F$-test score, the corresponding $p$ value computed is $p\simeq 0$, thereby signifying that the ANN model does not provide any statistically significant improvement over the $\Lambda$CDM estimates.
@@
Figure 4: Neural network reconstruction for distance modulus, $\mu(z)$ (black) with the standard deviation of the predictions plus the modeled error (yellow region) using LSST data (green dots). Left: With $\Lambda$CDM and CPL using the values of the parameter estimation using LSST data. Right: In comparison with data points from other SNIa surveys, Pantheon and Pantheon+ using $\Lambda$CDM model only.
@@
Table 2: Mean squared error between $\mu(z)$ predictions for (first row) three different models ($\Lambda$CDM, CPL and ANN) with the observed $\mu(z)$ from the LSST SNIa sim dataset and (second row) between the ANN reconstructed distance estimates with the corresponding theoretical estimates from $\Lambda$CDM and CPL.
@@ VII Conclusions
In this paper, we have provided a proof of concept, to forecast a deep learning-based non-parametric dark energy reconstruction model. Which uses the state of the art deep learning tools to find solutions to the open question of finding the most accurate dark energy model. While trying to maximize the benefits one can draw from the huge SNIa datasets that the LSST, is going to provide in a couple of years. This paper gives a flavour of how a deep learning based dark energy reconstruction model, for SNIa observations, can be ideally implemented with the actual observation of SNIa in the years to come, not only for the LSST but also from other similar programs (with suitable modifications). In this paper, we show unequivocally that our reconstructed ANN-based model does perform (slightly) better than the traditional $\Lambda$CDM and CPL models when tracing the observational data. But it also shows possible areas for future studies. For example, the increased standard deviation observed in the ANN model compared to the theoretical model (Fig. 5) underscores the need for further investigation. This work not only demonstrates the viability of deep learning for cosmological reconstruction analysis but also opens avenues for future research to refine these models for higher precision in understanding dark energy better.

### v3

@@ IV.2 Pantheon and Pantheon+ compilations
The Pantheon datasets are all publicly available in the form of unbinned Hubble diagram data. Pantheon has 1048 data points scolnic2018complete within redshifts between $z=0.01$ and $z=2.3$, and Pantheon+ has 1550 Type Ia between $z=0.001$ and $z=2.26$ brout2022pantheon+ . They are the latest compilations of the SNIa observations, expanding from the previously concluded confirmed Pan-STARRS1 survey 2002SPIE.4836..154K . Unlike the 5785 LSST simulated SNeIa, all of Pantheon’s data are spectroscopic in nature and it has slight higher redshift range than the redshift range that the LSST is forecast to observe.
@@ V Methodology
This section describes how we use the ANN to generate neural network reconstructions for the distance modulus. We employ the methodology presented in gomez2023neuralrecs using the hyperparameter tuning suggested in gomez2023neuralgenetic . We normalize the dataset, and split the 5785 LSST simulated SNeIa in 80% for training and 20% for validation to monitor possible under or overfitting. The loss function used is Mean Squared Error (MSE) and we train our neural networks through 2000 epochs. In our case, we use the TensorFlow library to implement our neural network model. Below we outline the steps:
@@
  • We perform an exploratory analysis of the dataset and then prepare the dataset for training our neural network. We consider the redshift ($z$) as input, and the distance modulus ($\mu$) with the error ($\sigma(\mu)$) conformed by the sum of their statistical and systematic errors.
  • We train different architectures of neural networks with Monte Carlo Dropout gal2015dropout (as in gomez2023neuralrecs ) to gain empirical knowledge about the range of hyperparameters. The Monte Carlo Dropout method randomly deactivates neurons during predictions after the neural network has been trained. This approach allows for estimating the uncertainties in the neural network’s predictions, is mathematically equivalent to approximating variational inference in deep Gaussian processes, and helps create more robust models by effectively reducing the risk of overfitting. We use the Monte Carlo dropout implementation from AstroANN astronnleung2019deep ; astronn . The use of Monte Carlo Dropout introduces probabilistic behavior in neural networks and aligns with the Bayesian neural networks paradigm gal2015dropout ; benatan2023enhancing . Predictions generated using this method should be represented by the mean and standard deviation of multiple predictions, each with different neurons turned off.
@@
  • To find a proper combination of hyperparameters, we set the search space for the use of nnogada (gomez2023neuralgenetic, ) and their implemented canonical genetic algorithms with elitism from the DEAP library DEAP_JMLR2012 ; de2014deap (for details on genetic algorithms we recommend Ref. medel2023cosmological ). In our case, we consolidated the search space with the number of layers $\in[3,4]$, the number of neurons for hidden layers $\in[100,200]$, and batch size $\in[8,16,32,64]$.
@@
  • The best combination found was 4 layers with 200 neurons in each one and a batch size of 32 with the MSE loss function value of 0.02528 (see Fig. 1).
  • We use the best configuration to build our artificial neural network and train it with LSST simulated data.
@@
  • Once the neural network is trained, we perform several predictions of the distance modulus using different redshift values, within the range of the training data, to generate a non-parametric reconstruction based on the LSST simulated SNIa data.
@@
  • To have an insight into the cosmological models considered in this work constrained by the LSST data, we perform an initial Bayesian inference using nested sampling with the dynesty library speagle2020dynesty and the SimpleMC code aubourg2015 ; simplemc . We also do it with Pantheon and Pantheon+, to compare the results obtained with the LSST simulations. This provides us with a baseline reference, before the subsequent ANN analysis. We compare the theoretical predictions and the reconstruction with the data, and we measure with the MSE metric to test which cases are in more agreement with the dataset. The values of the dark energy model parameters are previously fitted with Bayesian inference using the same data used for the neural reconstruction.
Given that the available LSST data is simulated and, for simplicity, we are using statistical errors of the distance modulus with a diagonal covariance matrix, we acknowledge that approaches such as those in gomez2023neuralrecs consider modeling non-diagonal covariance matrices; however, this is beyond the scope of our paper. During training, the ANN learns to predict the distance modulus along with its statistical errors given a redshift value from the dataset. Consequently, predictions are provided for each redshift corresponding to the pair (distance modulus, error). To generate a point in the reconstruction, we employ Monte Carlo Dropout, which uses the trained neural network to make multiple predictions by randomly deactivating different neurons. We then calculate the mean and standard deviation of these predictions, with the latter representing the intrinsic error from the ANN. This intrinsic error is added to the modeled statistical error.
We use the mean squared error (MSE) to compare the similarities between the neural reconstruction predictions, the theoretical models, and the SNIa compilations, as it offers a straightforward method for this analysis. Although other approaches, such as calculating Bayesian evidence or information criteria, are available for model comparison, we do not employ them due to the complexity and large number of parameters in neural networks. Our focus is on determining which theoretical model is most similar to the neural reconstructions, rather than seeking to develop a model superior to the theoretical ones.
For interested readers, the GitHub repository [^7] https://github.com/igomezv/LSST_DE_neural_reconstruction is available, with the data and the Python notebooks in which the hyperparameter tuning and reconstructions were performed.
@@ VI.1 Bayesian Analysis
As we mentioned in Section V, we perform a Bayesian parameter estimation of the $\Lambda$CDM (constrained to be flat) and CPL models using our LSST simulated data, Pantheon and Pantheon+ surveys. Table 1 includes the ranges of the priors used for the free parameters of both dark energy models. In addition, it shows the results of posterior distribution sampling and the Bayes factor to compare models. According to Fig. 2 and Table 1, we can notice that for $\Lambda$CDM the $\Omega_{m}$ is considerably better constrained with the LSST data than the other two SNIa compilations; while the $h$ parameter is practically not affected between the three different datasets. On the other hand, for the CPL model, the parameter estimation with LSST data is consistent with Pantheon and Pantheon+, with a better constraint for the $w_{0}$ parameter.
In all the cases the $-2\log\mathcal{L}_{max}$, equivalent to the chi-squared, (Table 1) obtains a lower value for the LSST dataset, which implies that this data obtains a tiny better fit for the cosmological models. In addition, something interesting is that using the LSST data, the Bayesian model comparison suggests that $\Lambda$CDM increases its advantage over CPL because according to Jeffrey’s scale, both Pantheon and Pantheon+ $\Lambda$CDM have moderate evidence with Bayes factors around 1.5, and for LSST data this scale is also moderate, but with a Bayes factor around 2.7, close to the boundary between moderate and strong. Please consider that we are using natural logarithm for Likelihood, Bayesian evidence, and Bayesian factor. This behavior can also be analyzed in Fig. 3, where the EoS for CPL using LSST data puts the $w=-1$ line for $\Lambda$CDM almost out of $1-\sigma$, while for Pantheon and Pantheon+, the most probable value for $w$ is statistically closer to the $\Lambda$CDM prediction. However, considering only the values of $-2\log\mathcal{L}_{max}$, for the three datasets, CPL has always a better fit to the data than $\Lambda$CDM; therefore, the Bayesian evidence could be very sensitive to the extra-parameters of CPL and, in future work, it could be worth to incorporate other model comparison techniques because Bayesian evidence has received some criticisms in cosmological data analysis efstathiou2008limitations ; nesseris2004comparison ; koo2022bayesian . However, our results clearly show that the LSST data can give more advantages while investigating for model selection tests in the future (more so with LSST $10$ years SNIa data) though in the current analysis our results from Table 1, are insufficient to decide in favor of any model.
@@
Evidence - - Moderate - Moderate - Moderate
Table 1: Bayesian analysis to $\Lambda$CDM and CPL models using Pantheon, Pantheon+ and LSST data. We can notice the priors of their parameters, their parameter estimation, $-2\log\mathcal{L}_{max}$, the log-Bayesian evidence $\log Z$, the log-Bayes factor $\log B$ of CPL and $\Lambda$CDM for each dataset. Log-Bayes factor is defined as $\log B_{ab}\equiv\log Z_{a}-\log Z_{b}$, and according to Jeffrey’s scale vazquez2012bayesian we can consider weak evidence if $10\leq|\Delta\log Z|<$ 1, moderate evidence if $31\leq|\Delta\log Z|<$ 3, strong evidence if $53\leq|\Delta\log Z|<$ 5, and decisive evidence if $|\Delta\log\mathcal{Z}|\geq 5$, in favour of the preferred model; if $\Delta\log Z$ is negative it is in favour of the second term, and if it is positive the first term is favoured hee2016bayesian . In our results, $\Lambda$CDM is always the preferred model.
@@
Figure 2: One and two-dimensional posterior distributions from Bayesian sampling for the free parameters of $\Lambda$CDM (left) and CPL (right) using SNIa data from Pantheon (1048 SNIa), Pantheon+ (1550 SNIa), and LSST simulations (5785 SNIa). For reference, dashed lines indicate the theoretical $\Lambda$CDM values of $w_{0}=-1$ and $w_{a}=0$.
@@ VI.2 Neural Reconstruction
Finally, after careful training of a neural network model with the LSST simulated data, we generated the model-independent reconstruction for $\mu(z)$ as can be seen in Fig. 4. In Fig.4 we plot the distance modulus ($\mu$) from the datasets of LSST sim and Pantheon, Pantheon+, and then we overlay with the reconstructed distance modulus (black dashed lines) along with their associated uncertainties from the ANN predictions plus the modeled errors (yellow regions). We also plot the theoretically computed $\mu(z)$, obtained from the Bayesian parameter estimation using the same data, from the two cosmological dark energy models of $\Lambda$CDM and CPL. From the left panel of Fig. 4 it can be seen that the reconstructed data points from the ANN model and the LSST observations (green points) are very close. Next, we computed the MSE for each of these three cases with the observed LSST dataset, shown in Table 2. From Table 2 we see that MSE for ANN is slightly better than $\Lambda$CDM or CPL. This implies that the artificial neural network (ANN) based non-parametric reconstruction model can indeed more closely replicate the data than the theoretical models, albeit by a small margin. Overfitting in the neural reconstruction is avoided by Monte Carlo Dropout, which works as a regularisation method, and with careful tuning of hyperparameters with genetic algorithms. The interesting thing about the neural model is that, although it is a data-driven model without any physical or cosmological assumptions, it is similar to LCDM and CPL.
The second row in Table 2 gives additional insight by comparing the MSE between the ANN reconstructed $\mu(z)$ and the corresponding $\Lambda$CDM or CPL based distances. It is worth noting that, due to the hyperparameter tuning of the ANN and the Monte Carlo dropout, we have a good model based on the data, which looks visually smooth and does not have any overfitting or underfitting in machine learning terms.
@@
In Fig. 5 we present residual plots for the distance modulus $\mu(z)$, computed as $\delta\mu\mid_{i}=\mu_{i}-\mu_{model}(z_{i})$ (where $\mu_{i}$ corresponds to the $i^{th}$ observation from the LSST sim data) for two cases, a) LSST sim data - $\Lambda$CDM theory and b) LSST sim data - ANN model reconstructed $\mu(z)$ (we compare the $\mu(z)$ residuals between the ANN model and the $\Lambda$CDM model only, as $\Lambda$CDM is seen to have bigger differences with the ANN reconstruction as seen in Table 2). It can be seen that there is minor difference between the residuals from $\Lambda$CDM and ANN model. However, we can notice the sensitivity of the ANN model from outliers, which can make the error margins get bigger. We performed an additional $F$-test to see if there is any statistically significant difference between the distance modulus residuals from $\Lambda$CDM and ANN model  [^8] The $F$-test is used to compare the variances of two independent samples to test the hypothesis that they are equal. It is defined as the ratio of the variances, $F=\frac{\sigma_{1}^{2}}{\sigma_{2}^{2}}$, where $\sigma_{1}^{2}$ and $\sigma_{2}^{2}$ are the sample variances. This statistic follows the $F$-distribution under the null hypothesis that the variances are equal snedecor1989statistical . From the $F$-test score, the corresponding $p$ value computed is $p\simeq 0$, thereby signifying that the ANN model does not provide any statistically significant improvement over the $\Lambda$CDM estimates.
@@
Figure 4: Neural network reconstruction for distance modulus, $\mu(z)$ (black) with the standard deviation of the Monte Carlo Dropout predictions plus the modeled error (yellow region) using LSST data (green dots). Left: With $\Lambda$CDM and CPL using the values of the parameter estimation using LSST data. Right: In comparison with data points from other SNIa surveys, Pantheon and Pantheon+ using $\Lambda$CDM model only.
@@
Table 2: Mean squared error between $\mu(z)$ predictions for (first row) three different models ($\Lambda$CDM, CPL and ANN) with the observed $\mu(z)$ from the LSST SNIa sim dataset and (second row) between the ANN reconstructed, or predicted, distance with the corresponding theoretical estimates from $\Lambda$CDM and CPL, the differences of the second row are useful to analyze what cosmological model is in more agreement with the data-based neural model.
@@ VII Conclusions
In this paper, we have provided a proof of concept, to test a deep learning-based non-parametric dark energy reconstruction model using SNIa simulated data from LSST. Which uses the state of the art deep learning tools to find solutions to the open question of finding the most accurate dark energy model. While trying to maximize the benefits one can draw from the huge SNIa datasets that the LSST, is going to provide in a couple of years. In this paper, we show that our reconstructed ANN-based model does perform slightly better than the traditional $\Lambda$CDM and CPL models when tracing the observational data. But it also shows possible areas for future studies. For example, the increased standard deviation observed in the ANN model compared to the theoretical model (Fig. 5) underscores the need for further investigation. This work not only demonstrates the viability of deep learning for cosmological reconstruction analysis but also opens avenues for future research to refine these models for higher precision in understanding dark energy better.
@@
VIII Data Availability
The data used in this work is available from authors upon request. The analysis notebooks are available in Github.